### v1

# FineCops-Ref: A new Dataset and Task for Fine-Grained Compositional Referring Expression Comprehension

Abstract
Referring Expression Comprehension (REC) is a crucial cross-modal task that objectively evaluates the capabilities of language understanding, image comprehension, and language-to-image grounding. Consequently, it serves as an ideal testing ground for Multi-modal Large Language Models (MLLMs). In pursuit of this goal, we have established a new REC dataset characterized by two key features: Firstly, it is designed with controllable varying levels of difficulty, necessitating multi-level fine-grained reasoning across object categories, attributes, and multi-hop relationships. Secondly, it includes negative text and images created through fine-grained editing and generation based on existing data, thereby testing the model’s ability to correctly reject scenarios where the target object is not visible in the image—an essential aspect often overlooked in existing datasets and approaches. Utilizing this high-quality dataset, we conducted comprehensive evaluations of both state-of-the-art specialist models and MLLMs. Our findings indicate that there remains a significant gap in achieving satisfactory grounding performance. We anticipate that our dataset will inspire new approaches to enhance visual reasoning and develop more advanced cross-modal interaction strategies, ultimately unlocking the full potential of MLLMs. Our code and the datasets are available at https://github.com/liujunzhuo/FineCops-Ref.
FineCops-Ref: A new Dataset and Task for Fine-Grained Compositional Referring Expression Comprehension
Junzhuo Liu1, Xuzheng Yang1, Weiwei Li1, Peng Wang1thanks: Corresponding author. 1University of Electronic Science and Technology of China junzhuo.cs@gmail.com, yangxuzheng@std.uestc.edu.cn, davelee.uestc@gmail.com, wangpeng8619@gmail.com
1 Introduction
Despite significant advancements in multimodal large language models (MLLMs), a critical challenge remains in ensuring these models’ responses are grounded in visual content rather than solely derived from linguistic cues Tong et al. (2024); Zhai et al. (2023); Miyai et al. (2024). Vision-language models often treat language as a bag of words, lacking meaningful engagement with word order, attributes, or relationships Ma et al. (2023); Thrush et al. (2022); Tong et al. (2024); Yuksekgonul et al. (2022), and exhibit poor grounding and spatial reasoning abilities Chen et al. (2024a); Tong et al. (2024); Zhang et al. (2024).
Current evaluation methods utilize Visual Question Answering or Image-Text Retrieval to evaluate the compositional reasoning or grounding abilities of MLLMs. However, these methods provide an indirect assessment of the models’ visual grounding capabilities. In contrast, the Referring Expression Comprehension (REC) task requires the model to directly output the bounding box coordinates based on a given expression, serving as an ideal testing ground for MLLMs.
Recent MLLMs, leveraging substantial grounding data Chen et al. (2023); Wang et al. (2023b, a) and specifically designed visual modules You et al. (2024); Li et al. (2024a), have achieved impressive results on common REC benchmarks like RefCOCO/+/g Yu et al. (2016). However, these benchmarks lack considerations of compositional reasoning, allowing models to perform well without understanding linguistic structure or even without the expression Cirik et al. (2018); Akula et al. (2020). Additionally, current vision-language models struggle with negative samples, where the target object is absent from the image Chen et al. (2020); Kurita et al. (2023); You et al. (2024). This limitation is further exacerbated by the lack of robustness in existing datasets, which fail to provide the necessary complexity and variability to thoroughly evaluate MLLMs.
In response, we introduce FineCops-Ref, a benchmark specifically designed to address these limitations. Our dataset introduces controlled difficulty levels, compelling MLLMs to perform fine-grained reasoning across object categories, attributes, and multi-hop relationships. We classify the difficulty levels based on the number of attributes and relationships necessary for locating the target object. For instance, if there is only one possible target in the image, the difficulty level is 1 regardless the complexity of the expression. If the model needs to understand at least two or more relationships and attribute information, the difficulty level is 3. Moreover, FineCops-Ref incorporates negative samples crafted through meticulous editing, testing the models’ resilience against misalignments and hallucinations, thereby assessing their true visual grounding capabilities.
Our comprehensive evaluation with state-of-the-art models reveals a significant gap in grounding performance, highlighting the need for advanced visual reasoning strategies. We present several core findings in our study. Firstly, for simple REC tasks with a difficulty level 1, traditional vision-language models, despite their relatively smaller parameter sizes, maintained a significant advantage. Secondly, all models exhibited poorer performance at difficulty levels greater than 1, while MLLMs demonstrated stronger capabilities under these conditions. In terms of negative data, all models showed weak performance, even in the simplest scenarios where the image does not contain an object matching the category specified in the expression. Additionally, we observed a positive correlation between precision on positive samples and recall with negative samples, with traditional vision-language models and MLLMs displaying different tendencies.
To enhance the fine-grained compositional reasoning capabilities of existing models, we employed the same pipeline used to construct our benchmark to create a rich training dataset that includes both positive and negative samples. Fine-tuning on this training dataset significantly improved model performance, with further improvements observed on the RefCOCO/+/g dataset. We make FineCops-Ref and the code for our data generation pipeline publicly available under the CC BY 4.0 License.
2 Related Works
Referring expression comprehension. The REC methods can generally be divided into two categories based on whether or not it uses LLMs: specialist and MLLMs. Specialists typically extract text and image features separately and perform multi-stage fusion Liu et al. (2023c); Yan et al. (2023); Kamath et al. (2021). Their training tasks often include various object location tasks. Recently, Zhao et al. (2024a) achieved excellent results on two visual grounding (VG) benchmarks by leveraging hard negative samples in training.
On the other hand, MLLMs directly input the projected visual features into the LLM. Recent methods aim to enhance grounding capabilities in MLLMs through dataset construction with coordinate information and additional visual modules. Common methods for datasets include transforming traditional visual datasets into an instruction-following format using templates Li et al. (2024b); Pramanick et al. (2023); Wang et al. (2023b), correlating object coordinates with existing captions Peng et al. (2024); Qi et al. (2024), and using GPT to generate question-and-answer pairs based on images, object coordinates, and captions You et al. (2024). The All-Seeing Project Wang et al. (2024) has recently introduced a new dataset (AS-1B) using a scalable data engine that incorporates human feedback and efficient models in the loop.
In terms of visual modules, some methods integrate additional visual components, such as GLaMM Rasheed et al. (2024) and LLaVA-Grounding Zhang et al. (2023), while others extract regional features to use as additional inputs Ma et al. (2024); Shao et al. (2024); You et al. (2024); Li et al. (2024a).
Figure 1: The data construction pipeline of FineCops-Ref. Given an image, we first generate paths based on its scene graph. Then, we fill paths into templates and obtain the positive referring expression through LLM rewriting. Meanwhile, we utilize LLM to generate negative expressions, and based on this, we employ diffusion model to create fine-grained editing negative images.
Evaluation of Compositional Reasoning. Current multimodal models, including advanced MLLMs like GPT-4V, exhibit poor compositional reasoning, often treating language as a bag of words without considering word order, attributes, or relationships between objects Suhr et al. (2019); Ma et al. (2023); Diwan et al. (2022); Tong et al. (2024); Yuksekgonul et al. (2022). Common evaluation benchmarks involve constructing hard negative captions to test models’ capabilities, such as distinguishing between "a mug in some grass" and "some grass in a mug" Parcalabescu et al. (2022); Thrush et al. (2022); Ma et al. (2023). Hsieh et al. (2023) found that previous benchmarks have language biases and that a simple grammar model can distinguish negative captions. Some benchmarks focus on negative images Ray et al. (2023); Yarom et al. (2023); Zhang et al. (2024); Le et al. (2023), while others primarily focus on spatial relationships Zhang et al. (2024); Liu et al. (2023a); Yang et al. (2019); Chen et al. (2024a).
For REC tasks, Akula et al. (2020) critically examined RefCOCOg, showing that 83.7% of test instances do not require reasoning on linguistic structure, and proposed the Ref-Adv dataset, which perturbs original expressions to refer to different target objects.
CLEVR-Ref+ Liu et al. (2019) is a synthetic dataset emphasizing relationships, attributes, and linguistic logic. Cops-Ref Chen et al. (2020) and Ref-Reasoning Yang et al. (2020) use GQA scene graphs Hudson and Manning (2019) and rule-based methods to create large-scale compositional referring expression comprehension datasets in real-world scenarios. Cops-Ref additionally added distracting images based on attributes, relationships, and target names. GITM-MR Wu et al. (2023) explores mismatched relationship in the REC task. RefEgo Kurita et al. (2023) and OmniLabel Schulter et al. (2023) consider out-of-distribution scenarios where referred targets do not exist in the image.
This paper addresses the limitations of previous benchmarks by constructing a REC dataset that comprehensively evaluates the compositional understanding abilities of existing multimodal models.
3 FineCops-Ref
FineCops-Ref includes both positive and negative data. Figure 1 illustrates the data construction pipeline.
3.1 Creating Positive Data
Path Generation. We employ image scene graphs from GQA Hudson and Manning (2019) for path generation. The scene graphs contain detailed information about objects, attributes, and relations. To ensure accuracy, we first filter the objects based on their suitability as target or related objects. We leverage annotations from InstInpaint Yildirim et al. (2023) and applying additional filters such as keywords and object size.
Next, we generate several paths for each of the filtered objects, as show in Figure 1(a). To eliminate any ambiguity, we utilize unique attributes or relations to identify the target object that share the same category as other objects in the image. and we make sure that every generated path are unique.
Data categorization. We categorize positive expressions into three difficulty levels, depending on the complexity of fine-grained reasoning. Level 1 indicates that there are no objects in the image with the same category as the target object. In this case, model can locate the target without requiring contextual understanding. Level 2 signifies the presence of an object with the same name as the target in the image, and the target can be distinguished through one unique attribute or relation. Level 3 require at least two or more relationships and attribute information. The difficulty level is established based on the intricacy of fine-grained reasoning, rather than the complexity of the textual description.
Positive Negative
Benchmark Unconstrained Cops. Difficulty level Neg. text Neg. image Expression Expression Image
RefCOCO 10752 - -
RefCOCO+ 10615 - -
RefCOCOg 9,602 - -
Ref-reasoning 34,609 - -
Cops-ref 12586 - 37758
Ref-adv 9602 3704 -
Ours 9605 9814 8507
Table 1: Comparison between the proposed benchmark and other REC benchmarks. Unconstrained indicates the final expression is not constrained by the templates. Cops. indicates fine-grained compositional reasoning. On the right hand side, the test set count of each benchmark is listed.
Expression Generation. We first employ a data balancing technique that takes into account the proportion of each type, effectively minimizing bias in the scene graph. Subsequently, the generated paths are substituted into predefined templates to generate reference expressions. We detail the predefined templates in Appendix A.1.
To further augment the naturalness and diversity of these expressions, we leverage GPT-3.5-turbo to rewrite the referring expressions. By incorporating well-designed instruction and examples, we are able to achieve a more expansive range of linguistically varied and authentic expressions. Prompts used to rewrite are listd in Appendix A.4.
Human Filter. Owing to the inherent constraints of the scene graph annotation information, the data pertaining to levels 2 and 3 may contain inaccuracies, leading to non-uniqueness in the targets referenced. To address this, human annotators filtered this portion of the data manually. Details refer to Appendix A.5.
3.2 Generating Negative Data
To conduct a thorough and systematic assessment of the REC in existing MLLMs, we generate hard negatives from both textual and visual sources. Like positive data, negative data are categorized into different levels based on the difficulty. Level 1 signifies alterations made to the target object in the negative data, which are relatively straightforward for the model to identify. Level 2 involves modifications to the related objects, disrupting the contextual information and posing a greater challenge for existing models to recognize.
Generating Negative Expressions. Our array of negative expressions encompasses a wide range of challenging types. We draw inspiration from CREPE Ma et al. (2023) and SUGARCREPE Hsieh et al. (2023) to consider various forms of hard negatives. In total, FineCops-Ref covers 5 fine-grained types of hard negative expressions. These types can be broadly classified into two categories: replace and swap. Replace involves generating a negative expression by substituting a portion of the original expression, whether it is an object, an attribute, or a relation. During replacement, we tested the output quality of several approaches. Ultimately, we found that LLM Replace performed the best. For more information, please refer to Appendix A.3. We utilize LLM to determine the most appropriate negative word, ensuring that the negative expression is genuinely negative while only slightly deviating from the original expression. On the other hand, swap entails generating a negative expression by interchanging two attributes or objects within the same category. We further employ LLM to rewrite the new expression. Please refer to Appendix A.2 for more details.
Generating Negative Images. We consider the necessity of negative images from the following aspects. First, negative images enables a more thorough assessment of models’ visual parsing capabilities. Additionally, the evaluation conducted by Visualgptscore Lin et al. (2023) suggests that negative expressions may lack plausibility and fluency and can be detected by language prior.
We generate hard negative images that bears slight differences from the original, such as alterations in objects, attributes, or relations. When dealing with simple positional relationships, we employ horizontal flips. For more intricate modifications involving objects and attributes, we utilize PowerPaint Zhuang et al. (2023), an exceptional image inpainting model offering versatility, to perform precise edits on the image. To guide PowerPaint in editing the image, we utilize LLM-generated replacements as textual guides and the bounding box as a mask. Overall, FineCops-Ref encompasses 5 distinct types of challenging negative images. Please refer to Appendix A.2 for more details.
Negative Data Debiasing. During the generation of negative samples, it is inevitable that certain implausible and incoherent expressions, as well as unreasonable and easily distinguishable negative images, may emerge. To ensure the benchmark’s quality, we employed various techniques to filter out these unsuitable samples and further improve the quality of the benchmark. For negative expressions, we employ the Adversarial Refinement technique proposed by SUGARCREPE Hsieh et al. (2023). It helps mitigate biases and unintended artifacts in the dataset.
To ensure the exclusion of inappropriate and excessively unreasonable negative images, we employ a multi-step filtering process. First, we use CLIP Radford et al. (2021) to ensure that the similarity between the negative text and the positive image is lower than the similarity between the positive text and the positive image. Next, we apply a diffusion-generated inspection model, DIRE Wang et al. (2023c), to filter out excessively unnatural images, excluding those with scores exceeding 0.2. Subsequently, we use DINOv2 Oquab et al. (2023) to compute the image-image similarity between the positive and negative images, retaining the one with the highest DINOv2 score from the 10 candidate negative images.
Set Positive Negative expression Negative image
Train 163792 80451 -
Val 18455 9029 -
Test 9605 9814 8507
Table 2: Dataset statistics.
3.3 Statistics
FineCops-Ref consists of 9,605 positive expressions, 9,814 negative expressions, and 8,507 negative images. Table 1 provides a comparison between FineCops-Ref and other visual grounding benchmarks. FineCops-Ref combines the advantages of unconstrained expression, fine-grained compositional reasoning, difficulty level, and hard negatives at both textual and visual levels. Additionally, we partition the training set and validation set simultaneously as in Table 2. For more details, please refer to the Appendix A.2.
3.4 Metrics
To evaluate performance on positive data, we use the common metric Precision@k. When both positive and negative data are present in the test set, we treat the negative samples as distractors for the positive samples, and introduce two additional metrics:
Recall@k: We treat the REC task as a bounding box retrieval problem. For each negative sample paired with its corresponding positive sample, we first obtain the predicted bounding boxes from the model, along with their confidence scores for both positive and negative samples. These bounding boxes are then ranked based on their confidence scores. Recall@k measures the proportion of negative-positive pairs where at least one of the top $k$ predicted bounding boxes has an IoU greater than 0.5 with the ground truth bounding box. It specifically assesses the model’s ability to avoid assigning high confidence scores to negative samples. Formally, Recall@k is defined as:
[…]
where $N$ represents the total number of negative-positive pairs, and $\mathbbm{1}(\cdot)$ is an indicator function that equals 1 if the condition inside is true and 0 otherwise. The term $\text{IoU}_{i,j}$ refers to the overlap between the $j$-th predicted bounding box (ranked based on the confidence scores) and the ground truth bounding box for the $i$-th pair. Note that for negative samples, there is no ground truth box, meaning the IoU is 0.
Recall@k is commonly used in retrieval tasks to assess prediction accuracy in the presence of challenging negative samples. Ideally, the model should assign lower confidence scores to negative samples. In our study, we primarily report Recall@1. If the model consistently assigns lower confidence scores to negative samples compared to positive ones, Recall@1 should equal Precision@1.
AUROC: While Recall@k evaluates how well the model ranks individual negative samples relative to their corresponding positive samples, it does not offer a holistic view of confidence across the dataset. To address this, we use AUROC to measure the overall ability of the model to distinguish between positive and negative samples. AUROC measures the model’s ability to correctly rank positive samples higher than negative ones across the datasets, providing a holistic view of its discriminative power.
By combining Recall@k and AUROC, we ensure a comprehensive evaluation of the model’s performance in distinguishing between positive and negative samples in REC tasks. This dual approach addresses both specific ranking and overall confidence.
4 Experiment
Positive
Model L1 L2 L3 Avg.
Specialist
Mdetr 72.43 52.79 46.92 57.38
MM-GDINO-T 75.11 34.78 35.46 48.45
MM-GDINO-L 85.13 43.54 42.89 57.19
UNINEXT 59.95 43.60 40.98 48.18
[…]
MLLM
Shikra 64.64 50.29 43.95 52.96
Ferret-13B 68.24 54.88 47.56 56.89
GroundingGPT 71.01 53.35 49.89 58.08
Lenna 73.75 41.92 38.43 51.37
InternVL 51.40 45.07 43.92 46.80
CogVLM 74.59 62.49 57.11 64.73
CogCom 76.23 60.86 60.08 65.72
GPT4-V + SoM 55.94 45.94 49.29 50.39
[…]
Table 3: Evaluation results (Precision@1) on positive data. $\dagger$ indicates training with positive samples from the training set, and $\ddagger$ indicates training with the entire training set. The best results are in bold, and the second-best results are underlined. The same notation will be used in subsequent tables.
4.1 Evaluation settings.
We evaluates several representative models, including both traditional vision-language models (Specialist) and MLLMs. The models examined in this study include MDETR Kamath et al. (2021), MM-GDINO Zhao et al. (2024b); Liu et al. (2023c), UNINEXT Yan et al. (2023), Shikra Chen et al. (2023), Ferret You et al. (2023), GroundingGPT Li et al. (2024b), Lenna Wei et al. (2023), InternVL Chen et al. (2024b), CogVLM Wang et al. (2023a) and CogCom Qi et al. (2024). We use there open-source checkpoints to evaluate.
We additionaly evaluate the GPT4-VAchiam et al. (2023). Since GPT4-V’s ability to directly output bounding boxes is relatively limited, we use GPT4-V combined with the Set-of-Mark (SoM) Yang et al. (2023) to evaluate its performance. The Model source and implementation details are in Appendix B.
We also test the effectiveness of training with the training dataset constructed using our data generation pipeline. We fine-tuned MM-GDINO-T and CogVLM using the positive data from the constructed training set. In addition, we fine-tuned MM-GDINO-T using the entire training set. The training settings are in Appendix B.
We evaluate the models using Precision@1 for positive data; Recall@1 and AUROC for negative data. Specifically, models like MDETR and Lenna that have dedicated object detection modules can generate multiple detection boxes with associated confidence scores, allowing for direct computation of Recall@1 and AUROC. For models that generate coordinates as the text using an autoregressive approach, we use the probability of the coordinates tokens to calculate confidence Kurita et al. (2023); Mitchell et al. (2023).
REPLACE SWAP
Object Attribute Relation Object Attribute
Model L1 L2 L1 L2 L1 L2 L1 L2 L1 L2 Avg.
Specialist
MDETR 52.89 36.09 50.47 35.92 42.48 40.77 45.89 37.35 44.42 37.70 42.40
MM-GDINO-T 58.84 33.77 50.47 29.96 34.69 31.92 43.89 27.71 43.67 31.97 38.69
MM-GDINO-L 64.23 40.26 55.76 41.52 45.74 43.73 53.02 48.19 49.38 37.70 47.95
UNINEXT 47.83 33.70 44.66 34.30 39.51 35.61 45.31 37.35 41.69 31.97 39.19
[…]
MLLM
Shikra 44.99 33.11 41.25 33.03 35.78 39.85 42.27 39.16 39.70 32.79 38.19
Ferret-13B 38.38 33.01 37.57 34.48 35.58 34.69 38.69 34.94 35.73 35.25 35.83
GroundingGPT 42.24 35.13 40.14 33.75 37.51 36.72 41.77 39.76 35.24 39.34 38.16
Lenna 65.88 50.38 58.75 42.96 47.00 43.91 49.94 38.55 49.38 43.44 49.02
CogVLM 53.34 44.02 51.24 48.74 41.22 44.46 47.69 49.40 46.40 40.16 46.67
CogCom 57.96 44.91 54.65 44.04 45.81 41.70 51.03 43.98 47.39 36.89 46.84
[…]
Table 4: Evaluation results (Recall@1) on negative expressions.
REPLACE SWAP
Object Attribute Object Attribute Flip
Model L1 L2 L1 L2 L1 L1 L2 L1 L2 Avg.
Specialist
MDETR 58.15 42.85 51.70 37.95 48.86 49.49 44.76 44.29 42.22 46.70
MM-GDINO-T 58.46 40.73 44.75 37.61 46.25 51.33 28.67 39.50 40.94 43.14
MM-GDINO-L 66.35 49.45 54.93 49.05 55.05 62.63 46.85 45.21 46.48 52.89
UNINEXT 48.85 31.62 40.96 30.33 46.91 40.25 37.06 30.66 29.42 37.34
[…]
MLLM
Shikra 42.57 33.61 36.54 34.26 35.18 38.60 36.36 34.25 37.10 36.50
Ferret-13B 41.54 37.46 38.04 36.22 43.00 37.78 39.16 35.27 36.25 38.30
GroundingGPT 43.91 36.88 36.31 35.88 39.09 37.17 40.56 37.02 33.05 37.76
Lenna 66.88 51.19 54.38 39.34 47.56 49.08 43.36 33.98 30.92 46.30
CogVLM 51.11 49.01 43.49 46.10 50.49 53.80 49.65 43.74 37.74 47.24
CogCom 32.24 21.55 22.57 20.10 39.74 25.46 18.88 24.13 23.03 25.30
[…]
Table 5: Evaluation results (Recall@1) on negative images.
4.2 Evaluation on Positive data
The results shown in Table 3 indicate that categorizing the dataset by difficulty level is crucial, as the performance of the most of the models declines with increasing difficulty. Notably, for level 3, most models achieve a precision below 50%.
Specialist perform better on simple REC task. At level 1, models merely need to detect objects based on their names, aligning with the requirements of open-vocabulary object detection. It was observed that Grounding DINO, based on SWIN-L, achieved an accuracy of 85.13% under zero-shot settings. This leads to two conclusions. First, vision-language models focused on object detection exhibit strong capabilities in basic visual localization and object detection tasks, even in zero-shot scenarios, which is also supported by their superior performance on RefCOCO benchmark which mainly require the model to detect the obejct without consider the attribute and relation. Second, although multimodal large models excel in dialogue and language understanding, their basic object detection abilities still fall short of the standards required for truly general-purpose models.
MLLMs exhibit superior reasoning abilities. For levels 2 and 3, models need robust language comprehension due to the presence of many easily confusable objects in the images. However, most models do not demonstrate sufficient capability in this aspect.
Multimodal models based on large language models (LLMs) achieved better results in this regard, demonstrating that MLLMs possess stronger compositional reasoning abilities.
Figure 2: The relationship between Precision@1 (on positive samples) and Recall@1 (on positive and negative samples) for Specialist and MLLM models across different negative difficulty levels. Specialist models correlate strongly with easier negative samples (Negative level 1, PCC = 0.923), while MLLMs show a higher correlation with harder negatives (Negative level 2, PCC = 0.917), reflecting their differing focuses on compositional REC.
4.3 Evaluation on Negative data
The evaluation results for negative expressions and negative images are shown in Table 4 and Table 5, respectively. We can draw the following conclusions:
The models are highly sensitive to the specific locations of negative data. L1 and L2 represent the replacement of the target directly and the replacement of other parts of the expression, respectively. For most types of negative data, the recall for L1 is significantly higher than for L2. This indicates that most models can identify simple anomalies, such as changes in the main target or inconsistencies in relationships. However, for L2 negative data, all models perform poorly, further demonstrating that the models lack compositional reasoning abilities and do not pay attention to the complete structure of the sentences.
The models have poor understanding of relationships. Overall, the models show relatively good recognition capabilities for direct object replacements, where the target mentioned in the expression is entirely absent in the image. Their ability to recognize attributes is slightly weaker. The models struggle significantly with understanding relationships, including recognizing replaced relationships and altered word order, which aligns with findings from previous studies. An additional finding is that the models perform worse in recognizing negative data of the "swap attribute" type compared to direct attribute replacements, indicating limitations in the models’ ability to bind attributes accurately.
5 In depth analysis
5.1 What’s the relationship between Precision and Recall?
In Figure 2, we explored the relationship between Precision@1 and Recall@1 among models. It is clearly evident that Precision and Recall are positively correlated. This is consistent with the findings of Ma et al. (2023); Vaze et al. (2022), where the accuracy of models on positive samples typically correlates positively with their ability to identify or reject out-of-distribution (OOD) samples.
Additionally, we further analyzed the correlation of different model types with different levels of negatives. We discovered a particularly interesting phenomenon: the precision of Specialist models has a Pearson Correlation Coefficient (PCC) of 0.923 with Negative level 1, whereas the precision of MLLMs has a PCC of 0.917 with Negative level 2. This further confirms the differing tendencies of MLLM and Specialist models. Specifically, Specialist tend to learn the existence and attributes of targets, while MLLMs models focus more on compositional reasoning.
5.2 Is rewrite useful?
To verify the significance of rewriting benchmark data, we conducted comparative experiments where models were evaluated using both the original data and the rewritten data. As shown in Table 6, models achieved significantly better performance on the evaluation benchmark without rewriting.
For positive data, using template-generated data always places the subject at the beginning of the sentence and has a very clear linguistic structure, which does not adequately assess the model’s language understanding abilities. For negative data, without rewriting, there are issues with non-fluency and nonsensicalityHsieh et al. (2023), which can be easily detected by text-only models such as Grammar Morris et al. (2020) and Vera Liu et al. (2023b).
Model Rewrite Percision@1 Recall@1
MM-GDINO-T 50.23 44.42
MM-GDINO-T 48.45 38.69
CogVLM 71.18 52.34
CogVLM 64.73 46.67
Grammar - 54.63
Grammar - 50.21
Table 6: Ablation study on the effect of rewriting the benchmark dataset. The reported metrics are the average Precision@1 and Recall@1 scores.
RefCOCO RefCOCO+ RefCOCOg
Model val test-A test-B val test-A test-B val test
CogVLM 92.76 94.75 88.99 88.68 92.91 83.39 89.75 90.79
[…]
Table 7: Evaluation results (Precision@1) on RefCOCO/+/g. The results of CogVLM come from the original paper.
5.3 Evaluation on RefCOCO
We additionally validated the performance of the CogVLM fine-tuned on our training set with RefCOCO/+/g benchmarks. As shown in Table 7, our model outperformed the original CogVLM in all validation and test sets. This result demonstrates the high quality and generalization capabilities of our dataset.
6 Conclusion
In this work, we introduced FineCops-Ref, a novel dataset for fine-grained compositional referring expression comprehension with varying difficulty levels and negative samples. Our evaluations reveal that while current MLLMs perform well on traditional REC benchmarks, they struggle with advanced compositional reasoning and accurate rejection of negative samples. Our dataset provides strong support for the evaluation of the model’s compositional grounding ability, and the training set can also serve as a good supplement to existing training data for compositional REC. We hope FineCops-Ref can inspire further research into enhancing compositional visual grounding.
7 Limitations
We employ LLMs and diffusion models for data generation, which inevitably introduce some hallucinations. Despite manual filtering of the benchmark dataset, hallucinations still persist in the training set. Additionally, while the models fine-tuned on the proposed training set exhibit good performance, we still lack effective methods for effectively recognizing hallucinations and handling negative samples.
Furthermore, although REC can evaluate the grounding ability of the model, the relationship between performance on REC tasks and other tasks such as VQA still needs to be explored. We also lack a complete evaluation of the model’s conversational abilities, like grounded image captions.
Appendix A Dataset details
A.1 Predefined templates
We have meticulously crafted a variety of templates tailored to suit different sentence structures, encompassing a range of 1-3 templates per structure. Examples of templates and corresponding expressions are shown in Table 8.
Type Exemplar templates Expression examples
0_hop The <$att_{0}$> <$obj_{0}$>. The white plate.
1_hop The <$att_{0}$> <$obj_{0}$> is <$rel_{0}$> the <$att_{1}$> <$obj_{1}$>. The giraffe is to the right of the trees.
and The <$att_{0}$> <$obj_{0}$> <$rel_{0}$> the <$att_{1}$> <$obj_{1}$> and <$rel_{1}$> the <$att_{2}$> <$obj_{2}$>. The balding man wearing the green shirt and to the left of the green trees.
2_hop The <$att_{0}$> <$obj_{0}$> is <$rel_{0}$> the <$att_{1}$> <$obj_{1}$> that is <$rel_{1}$> <$att_{2}$> <$obj_{2}$>. The blue, colorful and running train is on the bridge that is behind the green tree.
same_attr The <$obj_{0}$> sharing the <$rel_{0}$> as the <$obj_{1}$>. The plate that has the same color as the rice.
same_attr_2hop The <$obj_{0}$> that has the <$rel_{0}$> as the <$obj_{1}$> that <$rel_{1}$> the <$att_{0}$> <$obj_{2}$>. The table sharing the same color as the towels that to the right of the robe.
Table 8: Examples of expression type. $obj_{0}$ denotes the target object, while $obj_{1,2}$ denote the related objects. $att_{0,1,2}$ and $rel_{0,1}$ denote the corresponding attributes and relations, respectively.
Set L1 L2 L3 Sum.
Train 134466 25282 4044 163792
Test 5730 3404 471 9605
Val 15126 2884 445 18455
Table 9: Positive expressions Statistics. FineCops-Ref covers 3 difficult levels of positive expressions, split into train/test/val.
REPLACE SWAP
Set Object Attribute Relation Object Attribute Sum.
Train 29287 20678 14825 10062 5599 80451
Test 3951 1725 1891 1722 525 9814
Val 3308 2344 1676 1070 631 9029
Table 10: Hard negative expressions Statistics. FineCops-Ref covers 5 fine-grained types of hard negative expressions, split into train/test/val.
REPLACE SWAP
Set Object Attribute Object Attribute Flip Sum.
Test 4171 1844 307 630 1555 8507
Table 11: Hard negative images Statistics. FineCops-Ref covers 5 fine-grained types of hard negative images.
A.2 Examples of dataset
Difficulty levels. We categorize positive expressions into three levels, depending on the complexity of fine-grained reasoning. The difficulty criterion is established based on the intricacy of fine-grained reasoning, rather than the complexity of the textual description. Figure 3 showcases exemplary data ranging in difficulty levels.
Figure 3: Positive expressions of different difficulty levels.
Syntactic structure types. Meanwhile, following the syntactic structure, we categorize regular expressions into six types. $obj_{0}$ represents the target object, while $obj_{1}$ and $obj_{2}$ represent the related objects. "0_hop" indicates that the expression only involves $obj_{0}$, "1_hop" indicates that the expression mentions both $obj_{0}$ and $obj_{1}$. "And" and "2_hop" encompass $obj_{0,1,2}$. In "and," $obj_{1}$ and $obj_{2}$ are in a coordinated relationship, whereas in "2_hop," they are in a progressive relationship. "Same_attr" and "same_attr_2hop" restrict the relationship between $obj_{0}$ and $obj_{1}$ to the same attribute. Figure 4 showcases exemplary data ranging in syntactic structure types.
Figure 4: Positive expressions of different syntactic structure types.
Negative images. Figure 5 illustrates negative images generated by different methods.
Figure 5: Negative images generated by different methods.
Dataset statistics. For positive expressions and negative expressions, we split the dataset into train, test, and val sets. Specifically, positive expressions are classified based on levels, as detailed in Table 9. Negative expressions are classified based on types, as detailed in Table 10. While for negative images, we only generated them in the test set, categorized by type. Refer to Table 11 for more details.
A.3 Method to generate negative expressions
During our exploration into generating negative expressions, we delved into various methods to enhance the process. These methods encompassed the following approaches:(1) Predefined replace list: This method involves utilizing a predefined list of replacement words to substitute specific words. Although simple, it suffers from limited diversity and substantial bias. (2) Bert fill-mask: Employing this technique involves masking the original word and employing Bert to fill in the replacement. However, this method proves to be unstable and does not guarantee that the original word and its replacement belong to the same category. (3) LLM replace: This approach prompts the Language Model to generate the replacement word. It offers a high degree of richness and delivers reasonable outputs. Nonetheless, it requires a significant amount of time. In Table 12, we compare the outputs of these three methods using the vera and grammar score. The results indicate that LLM replace emerges as the optimal choice, both grammatically and logically.
Method Vera Grammar
Predefined replace list 70 55
Bert fill-mask 57 40
LLM replace 61 50
Table 12: Vera and Grammar score of different method’s output. The closer the score is to 50, the higher the quality of the data.
A.4 Examples of prompt
Prompt to rewrite expressions. We encourage the LLM to rephrase the given statement, aiming for rich and organic expressions while ensuring consistency throughout. Context learning was employed to integrate manually rewritten examples into the prompt. Additionally, to address any potential hallucinations, the LLM was instructed to include the original expression once in the output. Furthermore, two additional expressions were generated to enhance the diversity of the results. Figure 6 illustrates the prompt provided to the LLM for the rewriting task.
Figure 6: Prompt used for rewriting expressions.
Prompt to generate negative expressions. We prompt LLM to replace the specified word in the expression. LLM is required to find a misleading word that falls within the same category as the original word, yet contradicts its meaning. To ensure consistent and precise results, we have implemented stringent guidelines within the prompt. Furthermore, we have employed context-based learning by incorporating manually replaced instances in the prompt. Figure 7 illustrates the prompt provided to LLM for finding misleading words.
Figure 7: Prompt used for generating REPLACE negative expressions.
A.5 Human filter
We use the following prompt to guide human annotators to filter data. Program used for human filter see Figure 8.
Please determine whether the natural language description can accurately and unambiguously refer to the subject target contained within the red box in the image. In the image, the red box marks the subject target, while the green and blue boxes represent other objects mentioned in the language description. Please follow the guidelines below:
1. Carefully consider the attributes and relationships in the natural language description to ensure they accurately correspond to the image; otherwise, select “Wrong expression,”
2. Confirm whether the natural language description can uniquely refer to the target contained within the red box. If there are multiple possible targets, select “Ambiguous,”
3. If the natural language description is difficult to understand or cannot correctly refer to the subject target, please select "Wrong expression.”
Figure 8: Program used for human filter.
Appendix B Implementation details
B.1 Hardware information
All experiments are run on a machine with an Intel(R) Xeon(R) Gold 6348 CPU with a 512G memory and four 80G NVIDIA RTX A800 GPUs.
B.2 Dataset sources
We obtain all existing datasets from their original sources released by the authors. We refer readers to these sources for the dataset licenses. To the best of our knowledge, the data we use does not contain personally identifiable information or offensive content.
B.3 Model configuration
Model sources. We detail the sources of the pretrained models we use in the paper.
B.4 Experiments details
Evaluation details. We obtain the bounding box coordinates and confidence scores predicted by the model on our benchmark, and then calculate the metrics.
  • Specialist: We use the official inference code to perform inference and record the bounding box coordinates and confidence scores of the output.
  • MLLMs: We use the official inference code for inference and record the bounding box coordinates. The confidence score is calculated using the sum of the log probabilities of the coordinate tokens Kurita et al. (2023); Mitchell et al. (2023).
  • GPT-4V+SoM: Following the SoM Yang et al. (2023), we first use MM-GDINO to obtain candidate bounding boxes. Then, we draw these bounding boxes and corresponding labels on the image and ask GPT-4v to choose the label. To save costs, testing was conducted on a sample of 5k instances.
Training detials. We detail the dataset and hyper-parameters used in training our own models.
  • MM-GDINO-T: We trained the model with a batch size of 32. The AdamW optimizer was used with a learning rate of 0.0002 and a weight decay of 0.0001. The learning rate was adjusted using a MultiStepLR scheduler. The training ran for 5 epochs. For negative samples, the ground truth bounding box was set as empty.
  • CogVLM: We followed the provided template and performed instruction tuning with the joined training set of ours and RefCOCO/+/g. The training was done with lora Hu et al. (2022) and a batch size of 32, using the AdamW optimizer with a learning rate of 0.0002 and a weight decay of 0.0001. The training ran for 1 epoch, with a cosine learning rate schedule.
Appendix C Detailed evaluation results
AUROC results. The experimental results of AUROC exhibit a similar trend to Recall, further confirming the following observations: (1) The models are highly sensitive to the specific locations of negative data. (2) The models have a poor understanding of relationships. Specifically, Lenna performs averagely on positive data but shows good performance on negative data. This suggests that Lenna possesses good discrimination ability but lacks visual localization capability.
REPLACE SWAP
Object Attribute Relation Object Attribute
Model L1 L2 L1 L2 L1 L2 L1 L2 L1 L2 Avg.
Specialist
MDETR 63.58 51.89 58.75 52.64 54.92 54.26 59.60 54.11 56.33 51.38 55.75
MM-GDINO-T 66.02 49.66 57.50 48.85 49.88 49.78 56.80 49.50 55.87 55.93 53.98
MM-GDINO-L 66.73 49.93 58.35 50.21 51.93 53.57 60.51 55.47 54.88 54.72 55.63
UNINEXT 61.24 51.39 57.59 51.62 54.35 52.22 58.57 52.05 57.07 49.58 54.57
[…]
MLLM
Shikra 58.57 51.14 55.37 52.96 52.67 52.88 57.07 51.44 55.04 48.42 53.56
Ferret-13B 52.44 49.34 49.39 48.80 50.17 48.24 51.07 48.80 49.93 50.04 49.82
GroundingGPT 55.14 50.90 50.76 49.45 50.04 48.15 53.11 49.44 49.83 50.51 50.73
Lenna 76.46 63.93 64.29 52.66 56.92 53.56 59.98 51.22 56.96 48.87 58.49
CogVLM 60.60 51.40 55.66 52.96 51.95 53.77 55.14 55.04 53.09 55.47 54.51
CogCom 63.47 52.51 56.83 52.60 53.28 51.83 58.60 54.08 54.87 49.79 54.79
[…]
Table 13: Evaluation results (AUROC) on negative expressions.
REPLACE SWAP
Object Attribute Object Attribute Flip
Model L1 L2 L1 L2 L1 L1 L2 L1 L2 Avg.
Specialist
MDETR 64.00 56.20 58.02 53.69 60.89 58.72 55.63 55.01 53.42 57.29
MM-GDINO-T 64.32 57.51 53.27 55.81 58.74 58.76 55.09 51.72 53.43 56.52
MM-GDINO-L 68.00 58.05 55.90 56.08 59.96 62.81 58.08 51.87 53.04 58.20
UNINEXT 62.13 53.92 54.80 52.44 63.20 57.49 50.76 51.14 49.57 55.05
[…]
MLLM
Shikra 55.94 50.40 50.92 51.36 52.47 56.64 47.08 51.57 51.45 51.98
Ferret-13B 56.09 52.61 51.01 51.20 55.78 53.80 49.49 51.24 50.99 52.47
GroundingGPT 56.75 50.86 48.52 49.37 54.09 51.76 50.67 52.84 47.46 51.37
Lenna 74.71 65.17 60.27 55.85 59.24 59.08 52.47 50.25 49.42 58.50
CogVLM 58.62 55.88 51.03 55.24 56.71 56.29 55.81 52.04 51.47 54.79
CogCom 37.91 33.34 31.45 29.67 47.38 34.57 31.32 33.44 31.56 34.52
[…]
Table 14: Evaluation results (AUROC) on negative images.